solv the formula for the specifed variable
A=1/2h(b1+b2) solve for b1

Answer by tutorcecilia(2152)

A=1/2h(b1+b2) solve for b1
(2/h)(A)=(2/h)(1/2h(b1+b2)
2A/h=b1+b2
2A/h-b2=b1+b2-b2
(2A/h)-b2=b1
.
Check by plugging the value of b1 back into the original equation